# Thresholded Lasso for high dimensional variable selection

### v2

@@ 4.1 Discussions
Choosing the set $A_{0}$ as in Proposition 4.1 is rather arbitrary; one could for example, consider the set of variables that are strictly above $\lambda\sigma/2$. Compared with the almost exact sparse recovery result in Theorem 1.1 [47], we have relaxed the restriction on $\beta_{\min}$: rather than requiring all non-zero entries to be large in absolute values, we only require those in a subset $A_{0}$ to be recovered to be large. In the statement of Lemma 4.3, we assume the knowledge of the error bounds on various norms of $\beta_{\text{\rm init}}-\beta$ and $h=\beta_{\text{\rm init}}-\beta_{T_{0}}$ implicitly (hence the name of “oracle”). In obtaining (66), we may substitute the bounds as derived in Lemma 3.4 in (61), as trivially, for $A_{0}\subseteq T_{0}$ and $h_{A_{0}}=\beta_{\text{\rm init},A_{0}}-\beta_{A_{0}}$,
@@
Lemma 4.3 explains our results in Theorem 1.1 [46] as well as the numerical results. We also introduce $\breve{s}_{0}$ so that the dependency of $t_{0}$ on the knowledge of $s_{0}$ is relaxed; in particular, it can be used to express a desirable level of sparsity for the model $I$ that one wishes to select. In general, we may assume that $\beta_{\min,A_{0}}\geq 2D_{0}\lambda\sigma\sqrt{s_{0}}$ so as to have a large range of effective thresholding parameters; cf. Section 3. Thus one can increase $t_{0}$ as $\beta_{\min,A_{0}}$ increases in order to reduce the number of false positives while not increasing the number of false negatives from the active set $A_{0}$. In this case, it is also possible to remove variables from $T_{0}^{c}$ entirely by increasing the threshold $t_{0}$ while strengthening the lower bound on $\beta_{\min,A_{0}}$ by a constant factor. For example, we may set $t_{0}>\left\lVert\beta_{\text{\rm init},T_{0}^{c}}\right\rVert_{\infty}$, rather than $\left\lVert\beta_{\text{\rm init},T_{01}^{c}}\right\rVert_{\infty}$ as in (69). Clearly, $\left\lVert\beta_{\text{\rm init},T_{0}^{c}}\right\rVert_{\infty}=\left\lVert h_{T_{1}}\right\rVert_{\infty}\leq\left\lVert h_{T_{1}}\right\rVert_{2}=O_{P}(\lambda\sigma\sqrt{s_{0}})$ and hence
$\beta_{\min,A_{0}}\geq\sqrt{2}\left\lVert h_{T_{01}}\right\rVert_{2}$$\geq$ $\left\lVert h_{T_{0}}\right\rVert_{2}+\left\lVert h_{T_{1}}\right\rVert_{2}\text{ ensures }$
@@
which is sufficient for (68) to hold so long as $t_{0}\asymp\left\lVert\beta_{\text{\rm init},T_{0}^{c}}\right\rVert_{\infty}\leq\left\lVert h_{T_{1}}\right\rVert_{2}$. Then following the same analysis as in Theorem 4.4, we can recover $A_{0}$ while eliminating variables from $T_{0}^{c}$; cf. (70). In general, when the strong signals are close to each other in their strength, then a small $\beta_{\min,A_{0}}$ implies that we are in a situation with low signal to noise ratio (low SNR); one needs to carefully tradeoff false positives with false negatives as shown in our numerical results in Section 5 and the supplementary Section K. Such results and their formal statements are omitted from the present paper.
@@ 4.2 Model specification
We now pause to briefly describe our experiment setup so as to further discuss variable selection in the context of Theorem 4.4. Moreover, we generate a class of models with different sparsity $s$ and $\ell_{1}$ norm on $\beta_{T_{0}^{c}}$ to shed lights on this connection with the work of [43].
Let $\left\lvert\beta_{j}\right\rvert$s be ordered as in (14). Let $S=\mathop{\text{\rm supp}\kern 0.86108pt}(\beta)$ and $A_{0},T_{0}$ be as defined in Proposition 4.1. Let $A_{0}=[a_{0}]$ and $T_{0}=[s_{0}]$. Then $T_{0}\setminus A_{0}=\{a_{0}+1,\ldots,s_{0}\}$. Let $\left\lvert S\right\rvert=s$. We divide $\beta$ into 4 components and write $\beta=\beta^{(11)}+\beta^{(12)}+\beta^{(2)}+\beta^{(0)}\in\mathbb{R}^{p}$. The first 3 components contain non-zero coordinates. See Figure 1 for an illustration of the model specification, where we set for some $C_{a}>1$, $0<C_{m}\leq 1$, $c_{t}\geq 0$, and $\lambda=\sqrt{2\log p/n}$,
$\beta_{j}^{(11)}$ $=$ ±$\pm C_{a}\lambda\sigma\cdot 1_{1\leq j\leq a_{0}},\quad\beta_{j}^{(12)}=\pm C_{m}\lambda\sigma\cdot 1_{a_{0}<j\leq s_{0}},$ (72)
$\beta_{j}^{(2)}$ $=$ ±$\pm c_{t}(\sigma/\sqrt{n})\cdot 1_{s_{0}<j\leq s}\;\;\;\text{ and}\;\;\beta_{j}^{(0)}=0\cdot 1_{s<j\leq p},$
where unspecified coordinates in each component are again set to 0. Hence $\left\lVert\beta^{(12)}\right\rVert_{2}=C_{m}\lambda\sigma\sqrt{s_{0}-a_{0}}$ in (72). Since $C_{a}>1$, we have by (13),
$\sum_{j\leq a_{0}}\min(\beta_{j}^{2},\lambda^{2}\sigma^{2})$ $=$ $a_{0}\lambda^{2}\sigma^{2},\text{ since }\left\lvert\beta_{j}\right\rvert>\lambda\sigma\text{ for}\;j\in A_{0}=[a_{0}],$ (73)
@@
$(\ell_{2})\quad\left\lVert\beta^{(2)}\right\rVert_{2}$ $=$ $\left\lVert\beta_{{T_{0}^{c}}}\right\rVert_{2}\approx\lambda\sigma\sqrt{(1-C_{m}^{2})(s_{0}-a_{0})},\;\text{and}$ (77)
$(\ell_{1})\quad\left\lVert\beta_{T_{0}^{c}}\right\rVert_{1}$ $=$ $\sqrt{\left\lvert T_{0}^{c}\cap S\right\rvert}\left\lVert\beta_{T_{0}^{c}}\right\rVert_{2}=\left\lVert\beta^{(2)}\right\rVert_{2}^{2}/[c_{t}\sigma/\sqrt{n}]$ (78)
@@ 4.3 Conditions in [43]
Without loss of generality, we order $\left\lvert\beta_{1}\right\rvert\geq\left\lvert\beta_{2}\right\rvert\geq\ldots\geq\left\lvert\beta_{p}\right\rvert$ as in (14). For the Lasso solution $\beta_{\text{\rm init}}$, it is assumed that $\lambda_{n}\geq 2\lambda\sigma\sqrt{(1+a)c^{*}}$ in (2) in Theorem 3 in [43]. Hence we assume $\lambda_{n}\asymp 2\lambda\sigma\sqrt{(1+a)c^{*}}$ throughout our discussion. Roughly speaking, we interpret $H_{q}=\{1,\ldots,q\}$ as the set of large coordinates in $\beta$ that one aims to recover in the settings of [43]. For convenience, we denote by $H_{0}:=[p]\setminus H_{q}=\{q+1,\ldots,p\}$ and $T_{1}^{*}\subset H_{0}$ the index set $\{j\in H_{0}:\left\lvert\beta_{j}\right\rvert\geq\lambda\sigma\}$. The sparsity assumption by [43] is set with the $\ell_{1}$ sparsity on $\beta_{H_{0}}$,
@@
Here the $\tilde{O}$ notation may hide some constants including $c^{*}>1$ and $q^{*}>4r_{1}^{2}q$, where $q^{*},c_{*},c^{*},q,r_{1}$ (and $r_{2}$) are all allowed to depend on $n$; cf. (82) to (84). This target set $H_{q}$ is slightly more restrictive than the set $A_{0}=\{1,\ldots,a_{0}\}$, since in $H_{0}$, one may still see large signals; cf. (112) and the proof of Theorem 3 [43]. To resolve this discrepancy and to properly interpret results in [43], we first extend $H_{q}$ by $T_{1}^{*}\subset H_{0}$ and denote this extended set by $L(q)$, where
$L(q)$ ::=:= $H_{q}\cup T_{1}^{*}=\{j:\left\lvert\beta_{j}\right\rvert\geq\lambda\sigma\},\;\left\lvert T_{1}^{*}\right\rvert\leq\eta_{1}/(\lambda\sigma)=\tilde{O}(2r_{1}^{2}q),$
@@
which correctly depend on $q^{*}$. Recall in our example in Section 4.2, $\left\lVert\beta_{A_{0}^{c}}\right\rVert_{2}\leq(s_{0}-a_{0})^{1/2}\lambda\sigma$, where $H_{q}\subset A_{0}$ and $T_{0}^{c}\subseteq([p]\setminus L(q))\subseteq A_{0}^{c}\subset H_{0}$,
@@
Notice that the tight upper bound in (87) (with matching lower bound in Section C.1) has an extra $\sqrt{\log p}$ factor compared to the $\ell_{1}$ condition in (80), where it is understood that $q^{*}\asymp s_{0}$ in order for the error $\left\lVert\beta_{\text{\rm init}}-\beta\right\rVert_{2}$ and $\left\lVert X(\beta_{\text{\rm init}}-\beta)\right\rVert_{2}$ in Theorem 3 [43] to match those in Theorem 2.4; cf. (107). In summary, although we impose sparsity in the $\ell_{0}$ sense, the actual lower bound on $\left\lVert\beta_{A_{0}^{c}}\right\rVert_{1}$,
@@
cf. (115) and (C.1). On the other hand, suppose $q^{*}\asymp\sqrt{\log p}(s_{0}-a_{0})$ is allowed, then this extra $\sqrt{\log p}$ factor in (88) will inevitably appear in the upper bounds derived in (86) and (85), resulting in worse $\ell_{2}$ loss while simultaneously, the bounded (upper and lower) sparse eigenvalue conditions need to hold for design matrix $X$ with rank $q^{*}=\Omega((s_{0}-a_{0})\sqrt{\log p})$.
@@
$\ell_{1}$ error on $h$ versus on $\delta$. Recall $h=\beta_{\text{\rm init}}-\beta_{T_{0}}$ and $\delta=\beta_{\text{\rm init}}-\beta$. Since $\mathop{\text{\rm supp}\kern 0.86108pt}(\beta_{\text{\rm init}})\asymp q^{*}$ by (105), these results on the $\ell_{1}$ error $\left\lVert\delta\right\rVert_{1}$, where $\delta=\beta_{\text{\rm init}}-\beta$, and the Lasso support in [43], cf. (86) and (105), are different from Theorem 2.4 in the present paper, as we do not aim to bound the Lasso support directly. A more subtle point is that since $q^{*}\asymp\left|A_{1}\right|$, we can crudely interpret $q^{*}$ as the size of $\mathop{\text{\rm supp}\kern 0.86108pt}(\beta_{\text{\rm init}})$, given $q\ll q^{*}$;
@@
In our theory, indeed, the Lasso support can be much larger than $s_{0}$ while the $\ell_{1},\ell_{2}$ norm bounds on $h$ are quite tight in the sense of Theorem 2.4. Hence the SRC assumptions (83) are somewhat similar and closely related to the sparse eigenvalue conditions in Theorem 2.4, where $\textnormal{{RE}}(s_{0},4,X)$ also holds with $K(s_{0},4)<\infty$. On the one hand, the RSC condition is set at a rank $q^{*}\asymp\left|\mathop{\text{\rm supp}\kern 0.86108pt}(\beta_{\text{\rm init}})\right|$, which may be significantly larger than $s_{0}$ in our model. On the other hand, the $\ell_{2}$ and $\ell_{1}$ error on $h$ in Theorem 2.4 depend on $s_{0}$, as well as $K^{2}(s_{0},4)$, $\Lambda_{\max}(s-s_{0})\asymp\Lambda_{\max}(\log p(s_{0}-a_{0})/c_{t}^{2})$, $\Lambda_{\max}(2s_{0})$, and $\Lambda_{\min}(2s_{0})$ (cf. (29)), without an explicit condition on (82).
[…]
@@
As shown in the bottom two plots in Figure 2, where we set $(p,n,s_{0},a_{0},\gamma,C_{a})=(2048,1600,50,30,0.7,1.706)$, we observe that $\left\lVert\delta\right\rVert_{1}$ dominates $\left\lVert h\right\rVert_{1}$ consistently across all $s\in\{130,370,511\}$ as the Lasso penalty $\lambda_{n}=f_{p}\lambda\sigma$ increases. In particular while upon rescaling, all curves corresponding to different values of $s$ align well for the $\ell_{2}$ error $\left\lVert\delta\right\rVert_{2}$, $\left\lVert h_{T_{0}}\right\rVert_{2}$ (right plot) and the $\ell_{1}$ norm $\left\lVert h\right\rVert_{1}$, $\left\lVert h_{T_{0}}\right\rVert_{1}$, and $\left\lVert h_{T_{0}^{c}}\right\rVert_{1}$ (left plot), the same is not true for $\left\lVert\delta_{T_{0}^{c}}\right\rVert_{1}$ and $\left\lVert\delta\right\rVert_{1}$. However, we have by the triangle inequality,
$\left\lvert\left\lVert\delta\right\rVert_{1}-\left\lVert h\right\rVert_{1}\right\rvert$ $=$ $\left\lvert\left\lVert\delta_{T_{0}^{c}}\right\rVert_{1}-\left\lVert h_{T_{0}^{c}}\right\rVert_{1}\right\rvert\leq\left\lVert\beta-\beta_{T_{0}}\right\rVert_{1}=\left\lVert\beta_{T_{0}^{c}}\right\rVert_{1};$ (89)
@@
We show the lower bound on $\left\lVert\beta_{T_{0}^{c}}\right\rVert_{1}$ in (115). Hence the error $\left\lVert\delta_{T_{0}^{c}}\right\rVert_{1}$ increases as $s$ increases (as $c_{t}$ decreases) but also bounded, as predicted by (77), (78) and (89). See Sections 5.1 and C for details.

### v3

@@ 4.1 Discussions
Theorem 4.4 is an immediate corollary of Theorem 2.1, Lemma 2.7, and Lemma 4.3, except that we now set $\breve{s}_{0}=s_{0}$ everywhere in Lemma 4.3, and assume having an upper estimate $\breve{D}_{1}$ (resp. $\breve{D}_{0}$) of $D_{1}$ (resp. $D_{0}$), so as not to depend on an “oracle” telling us an exact value. Given all other parameters held invariant, a larger $\beta_{\min,A_{0}}$ (e.g., a larger $C_{a}$ for $\beta$ in Figure 1) will result in a tighter bound on the bias $\left\lVert\beta_{{\mathcal{D}}}\right\rVert_{2}^{2}$ in view of (57). When $\beta_{\min,A_{0}}^{2}$ dominates the upper bound on the RHS of (57), then $A_{0}\cap I^{c}=\emptyset$, since the loss of a single variable from $A_{0}$ will already saturate this upper bound in the order of $O(\lambda^{2}\sigma^{2}s_{0})$. Hence when $\beta_{\min,A_{0}}$ is strong enough in the sense of (4.3), there will be no false negatives from the set $A_{0}$, leading to the stronger and tighter bounds in (60) and (61) that also control false positives and the bias. This is validated in our experiments in Section 5.
In the statement of Lemma 4.3, we assume the knowledge of the error bounds on various norms of $\beta_{\text{\rm init}}-\beta$ and $h=\beta_{\text{\rm init}}-\beta_{T_{0}}$ implicitly (hence the name of “oracle”). In obtaining (62), we may substitute the bounds as derived in Lemma 3.3 in (58), as trivially, for $A_{0}\subseteq T_{0}$ and $h_{A_{0}}=\beta_{\text{\rm init},A_{0}}-\beta_{A_{0}}$,
@@
Lemma 4.3 explains our results in Theorem 1.1 [46] as well as the numerical results. We also introduce $\breve{s}_{0}$ so that the dependency of $t_{0}$ on the knowledge of $s_{0}$ is relaxed; in particular, it can be used to express a desirable level of sparsity for the model $I$ that one wishes to select. In general, we may assume that $\beta_{\min,A_{0}}\geq 2D_{0}\lambda\sigma\sqrt{s_{0}}$ so as to have a large range of effective thresholding parameters; cf. Section 3. Thus one can increase $t_{0}$ as $\beta_{\min,A_{0}}$ increases in order to reduce the number of false positives while not increasing the number of false negatives from the active set $A_{0}$.
In this case, it is also possible to remove variables from $T_{0}^{c}$ entirely by increasing the threshold $t_{0}$ while strengthening the lower bound on $\beta_{\min,A_{0}}$ by a constant factor. For example, we may set $t_{0}>\left\lVert\beta_{\text{\rm init},T_{0}^{c}}\right\rVert_{\infty}$, rather than $\left\lVert\beta_{\text{\rm init},T_{01}^{c}}\right\rVert_{\infty}$ as in (66). Clearly, $\left\lVert\beta_{\text{\rm init},T_{0}^{c}}\right\rVert_{\infty}=\left\lVert h_{T_{1}}\right\rVert_{\infty}\leq\left\lVert h_{T_{1}}\right\rVert_{2}=O_{P}(\lambda\sigma\sqrt{s_{0}})$ and hence the condition $\beta_{\min,A_{0}}\geq\sqrt{2}\left\lVert h_{T_{01}}\right\rVert_{2}\geq\left\lVert h_{T_{0}}\right\rVert_{2}+\left\lVert h_{T_{1}}\right\rVert_{2}$ ensures
@@
which is sufficient for (65) to hold so long as $t_{0}\asymp\left\lVert\beta_{\text{\rm init},T_{0}^{c}}\right\rVert_{\infty}\leq\left\lVert h_{T_{1}}\right\rVert_{2}$. Then following the same analysis as in Theorem 4.4, we can recover $A_{0}$ while eliminating variables from $T_{0}^{c}$; cf. (67). In general, if the strong signals are close to each other in their strength, then a small $\beta_{\min,A_{0}}$ implies that we are in a situation with low signal to noise ratio (low SNR); one needs to carefully tradeoff false positives with false negatives as shown in our numerical results in Section 5 and the supplementary Section I. Such results and their formal statements are omitted from the present paper.
@@ 4.2 Model specification
We now pause to briefly describe our experiment setup so as to further discuss variable selection in the context of Theorem 4.4. Moreover, we generate a class of models with different sparsity $s$ and $\ell_{1}$ norm on $\beta_{T_{0}^{c}}$ to shed light on this connection with the work of [43]. Let $\left\lvert\beta_{j}\right\rvert$’s be ordered as in (14). Let $S=\mathop{\text{\rm supp}\kern 0.86108pt}(\beta)$ and $A_{0},T_{0}$ be as defined in Proposition 4.1. Let $A_{0}=[a_{0}]$ and $T_{0}=[s_{0}]$. Then $T_{0}\setminus A_{0}=\{a_{0}+1,\ldots,s_{0}\}$. Let $\left\lvert S\right\rvert=s$. See Figure 1 for an illustration of the model specification. We divide $\beta$ into 4 components and write $\beta=\beta^{(11)}+\beta^{(12)}+\beta^{(2)}+\beta^{(0)}\in\mathbb{R}^{p}$, where we set for some $C_{a}>1$, $0<C_{m}\leq 1$, $c_{t}\geq 0$,
$\beta_{j}^{(11)}$ $=$ $\pm C_{a}\lambda\sigma\cdot 1_{1\leq j\leq a_{0}},\quad\beta_{j}^{(12)}=\pm C_{m}\lambda\sigma\cdot 1_{a_{0}<j\leq s_{0}},$ (69)
$\beta_{j}^{(2)}$ $=$ $\pm c_{t}(\sigma/\sqrt{n})\cdot 1_{s_{0}<j\leq s}\;\;\;\text{ and}\;\;\beta_{j}^{(0)}=0\cdot 1_{s<j\leq p},$
where $\lambda=\sqrt{2\log p/n}$ and the unspecified coordinates in each component are again set to 0. The first 3 components contain non-zero coordinates. Hence $\left\lVert\beta^{(12)}\right\rVert_{2}=C_{m}\lambda\sigma\sqrt{s_{0}-a_{0}}$ in (69). Since $C_{a}>1$, we have by (13),
[…]
@@
$(\ell_{2})\quad\left\lVert\beta^{(2)}\right\rVert_{2}$ $=$ $\left\lVert\beta_{{T_{0}^{c}}}\right\rVert_{2}\approx\lambda\sigma\sqrt{(1-C_{m}^{2})(s_{0}-a_{0})},$ (74)
$\text{and}\;(\ell_{1})\quad\left\lVert\beta_{T_{0}^{c}}\right\rVert_{1}$ $=$ $\sqrt{\left\lvert T_{0}^{c}\cap S\right\rvert}\left\lVert\beta_{T_{0}^{c}}\right\rVert_{2}=\left\lVert\beta^{(2)}\right\rVert_{2}^{2}/[c_{t}\sigma/\sqrt{n}]$ (75)
@@
We now show that the upper bound in (75) is essentially tight. First, the bound in (74) is tight, since
$\geq$ $(s_{0}-a_{0})(1-C^{2}_{m})\lambda^{2}\sigma^{2}-\lambda^{2}\sigma^{2}.$
$\sum_{j\in A_{0}^{c}}\left\lvert\beta_{j}\right\rvert$ $=$ $\left\lVert\beta^{(2)}\right\rVert_{1}+C_{m}\lambda\sigma(s_{0}-a_{0})=\Omega(\lambda\sigma(s_{0}-a_{0}){\sqrt{(\log p)}}/{c_{t}}).$
Finally, (78) follows from (71) and the fact that $\left\lVert\beta^{(12)}\right\rVert_{2}^{2}=\sum_{a_{0}<j\leq s_{0}}\min(\beta_{j}^{2},\lambda^{2}\sigma^{2})=C^{2}_{m}\lambda^{2}\sigma^{2}(s_{0}-a_{0})$.
@@ 4.3 Conditions in [43]
Without loss of generality, we order $\left\lvert\beta_{1}\right\rvert\geq\left\lvert\beta_{2}\right\rvert\geq\ldots\geq\left\lvert\beta_{p}\right\rvert$ as in (14). For the Lasso solution $\beta_{\text{\rm init}}$, it is assumed that $\lambda_{n}\geq 2\lambda\sigma\sqrt{(1+a)c^{*}}$ in (2) in Theorem 3 in [43]. Hence we assume $\lambda_{n}\asymp 2\lambda\sigma\sqrt{(1+a)c^{*}}$ throughout our discussion. Roughly speaking, we interpret $H_{q}=\{1,\ldots,q\}$ as the set of large coordinates in $\beta$ that one aims to recover in the settings of [43]. For convenience, we denote by $H_{0}:=[p]\setminus H_{q}=\{q+1,\ldots,p\}$ and $T_{1}^{*}\subset H_{0}$ the index set $\{j\in H_{0}:\left\lvert\beta_{j}\right\rvert\geq\lambda\sigma\}$. The sparsity assumption by [43] is set with the $\ell_{1}$-sparsity on $\beta_{H_{0}}$,
@@
Here the $\tilde{O}$ notation may hide some constants including $c^{*}>1$ and $q^{*}>4r_{1}^{2}q$, where $q^{*},c_{*},c^{*},q,r_{1}$ (and $r_{2}$) are all allowed to depend on $n$; cf. (82) to (84). This target set $H_{q}$ is slightly more restrictive than the set $A_{0}=\{1,\ldots,a_{0}\}$, since in $H_{0}$, one may still see large signals; cf. the supplementary (113) and the proof of Theorem 3 [43]. To resolve this discrepancy and to properly interpret results in [43], we first extend $H_{q}$ by $T_{1}^{*}\subset H_{0}$ and denote this extended set by $L(q)$, where
$L(q)$ $:=$ $H_{q}\cup T_{1}^{*}=\{j:\left\lvert\beta_{j}\right\rvert\geq\lambda\sigma\},\;\left\lvert T_{1}^{*}\right\rvert\leq\eta_{1}/(\lambda\sigma)=\tilde{O}(2r_{1}^{2}q),$
@@
which correctly depend on $q^{*}$. Recall in our example in Section 4.2, $\left\lVert\beta_{A_{0}^{c}}\right\rVert_{2}\leq(s_{0}-a_{0})^{1/2}\lambda\sigma$, where $H_{q}\subset A_{0}$ and $T_{0}^{c}\subseteq([p]\setminus L(q))\subseteq A_{0}^{c}\subset H_{0}$, and
@@
Notice that the tight upper bound in (87) (with matching lower bound) has an extra $\sqrt{\log p}$ factor compared to the $\ell_{1}$-sparsity condition in (80), where it is understood that $q^{*}\asymp s_{0}$ in order for the error $\left\lVert\beta_{\text{\rm init}}-\beta\right\rVert_{2}$ and $\left\lVert X(\beta_{\text{\rm init}}-\beta)\right\rVert_{2}$ in Theorem 3 [43] to match those in Theorem 2.4; cf. the supplementary (108). In summary, although we impose sparsity in the $\ell_{0}$ sense, the actual lower bound on $\left\lVert\beta_{A_{0}^{c}}\right\rVert_{1}$,
@@
cf. (79) and (4.2). On the other hand, suppose $q^{*}\asymp\sqrt{\log p}(s_{0}-a_{0})$ is allowed, then this extra $\sqrt{\log p}$ factor in (88) will inevitably appear in the upper bounds derived in (86) and (85), resulting in worse $\ell_{2}$ loss while simultaneously, the bounded sparse eigenvalue conditions need to hold for design matrix $X$ with rank $q^{*}=\Omega((s_{0}-a_{0})\sqrt{\log p})$. See the supplementary Section 7.
@@
$\ell_{1}$ error on $h$ versus on $\delta$. Recall $h=\beta_{\text{\rm init}}-\beta^{\mathop{\natural}\kern 0.60275pt}({T_{0}})$ and $\delta=\beta_{\text{\rm init}}-\beta$. Since $\mathop{\text{\rm supp}\kern 0.86108pt}(\beta_{\text{\rm init}})\asymp q^{*}$ by the supplementary (106), these results on the $\ell_{1}$ error $\left\lVert\delta\right\rVert_{1}$, where $\delta=\beta_{\text{\rm init}}-\beta$, and the Lasso support in [43], cf. (86) and the supplementary (106), are different from Theorem 2.4 in the present paper, as we do not aim to bound the Lasso support directly. A more subtle point is that since $q^{*}\asymp\left|A_{1}\right|$, we can crudely interpret $q^{*}$ as the size of $\mathop{\text{\rm supp}\kern 0.86108pt}(\beta_{\text{\rm init}})$, given $q\ll q^{*}$;
@@
In our theory, indeed, the Lasso support can be much larger than $s_{0}$ while the $\ell_{1}$-, $\ell_{2}$-norm bounds on $h$ are quite tight in the sense of Theorem 2.4. Hence the SRC assumptions (83) are somewhat similar and closely related to the sparse eigenvalue conditions in Theorem 2.4, where $\textnormal{{RE}}(s_{0},4,X)$ also holds with $K(s_{0},4)<\infty$. On the one hand, the SRC condition is set at a rank $q^{*}\asymp\left|\mathop{\text{\rm supp}\kern 0.86108pt}(\beta_{\text{\rm init}})\right|$, which may be significantly larger than $s_{0}$ in our model. On the other hand, the $\ell_{2}$ and $\ell_{1}$ error on $h$ in Theorem 2.4 depend on $s_{0}$, as well as $K^{2}(s_{0},4)$, $\Lambda_{\max}(2s_{0})$, and $\Lambda_{\min}(2s_{0})$ (cf. (30)), and $\Lambda_{\max}(s-s_{0})\asymp\Lambda_{\max}(\log p(s_{0}-a_{0})/c_{t}^{2})$, without an explicit condition on (82).
We demonstrate the tightness of these theoretical bounds in Figures  2 and 3. Consider (69) in particular. In our experiments, $c_{t}$ is an absolute constant $\in[0.527,1.658]$ (See Table 2), and the signs of the non-zero values in $\beta^{(2)}$ are chosen from $\{\pm 1\}$ at random. Hence, we have a longer tail (in the sense of a larger $\left\lVert\beta^{(2)}\right\rVert_{1}$ with many small coordinates) and we expect the Lasso estimate to have a larger $\left\lVert\delta\right\rVert_{1}$. See Table 2, where we fix $\sigma=1$, $a_{0}=30$, $s_{0}=50$, and $C_{m}=1/\sqrt{2}$ for $s>s_{0}$.
@@
As shown in the bottom two plots in Figure 2, where we set $(p,n,s_{0},a_{0},\gamma,C_{a})=(2048,1600,50,30,0.7,1.706)$, we observe that $\left\lVert\delta\right\rVert_{1}$ dominates $\left\lVert h\right\rVert_{1}$ consistently across all $s\in\{130,370,511\}$ as the Lasso penalty $\lambda_{n}=f_{p}\lambda\sigma$ varies. In particular while upon rescaling, all curves corresponding to different values of $s$ align well for the $\ell_{2}$-error, namely, $\left\lVert\delta\right\rVert_{2}$, $\left\lVert h_{T_{0}}\right\rVert_{2}$ (right plot) and the $\ell_{1}$-norm $\left\lVert h\right\rVert_{1}$, $\left\lVert h_{T_{0}}\right\rVert_{1}$, and $\left\lVert h_{T_{0}^{c}}\right\rVert_{1}$ (left plot), the same is not true for $\left\lVert\delta_{T_{0}^{c}}\right\rVert_{1}$ and $\left\lVert\delta\right\rVert_{1}$. However, we have by the triangle inequality,
$\left\lvert\left\lVert\delta\right\rVert_{1}-\left\lVert h\right\rVert_{1}\right\rvert$ $=$ $\left\lvert\left\lVert\delta_{T_{0}^{c}}\right\rVert_{1}-\left\lVert h_{T_{0}^{c}}\right\rVert_{1}\right\rvert\leq\left\lVert\beta_{T_{0}^{c}}\right\rVert_{1}.$ (89)
@@
We show the lower bound on $\left\lVert\beta_{T_{0}^{c}}\right\rVert_{1}$ in (79). Hence the error $\left\lVert\delta_{T_{0}^{c}}\right\rVert_{1}$ increases as $s$ increases (as $c_{t}$ decreases) but also remains bounded, as predicted by (74), (75) and (89). See Sections 5.1.